### v2

# Towards Interpretable and Efficient Attention: Compressing All by Contracting a Few

Abstract
Attention mechanisms in Transformers have gained significant empirical success. Nonetheless, the optimization objectives underlying their forward pass are still unclear. Additionally, the quadratic complexity of self-attention is increasingly prohibitive. Unlike the prior work on addressing the interpretability or efficiency issue separately, we propose a unified optimization objective to alleviate both issues simultaneously. By unrolling the optimization over the objective, we derive an inherently interpretable and efficient attention mechanism, which compresses all tokens into low-dimensional structures by contracting a few representative tokens and then broadcasting the contractions back. This Contract-and-Broadcast Self-Attention (CBSA) mechanism can not only scale linearly but also generalize existing attention mechanisms as its special cases. Experiments further demonstrate comparable performance and even superior advantages of CBSA on several visual tasks. Code is available at this https URL.
1 Introduction
The Transformer architecture and attention mechanisms [1] have been widely applied across diverse areas, including computer vision [2, 3], natural language processing [4, 5], and scientific discovery [6]. Nonetheless, a series of puzzling phenomena have been uncovered in Transformers, such as the emergent segmentation properties [7], the in-context learning ability [8], the attention collapse [9, 10] and the extreme-token phenomena [11]. At the same time, the quadratic computational and memory complexity of self-attention with respect to the sequence length impedes its broader applications in real-time systems [12], as well as the processing of long documents [13] and high-resolution images [14].
In light of these challenges, it is crucial to demystify attention mechanisms and enhance the efficiency of self-attention. Over the past few years, remarkable advances have been made in addressing the interpretability or efficiency issue separately. On one hand, attention mechanisms can be interpreted as unrolling an optimization objective grounded in clustering [15], compression [16], or energy minimization [17]. These ante-hoc, inherently interpretable perspectives are generally considered more rigorous than post-hoc explanations [18]. On the other hand, numerous techniques have been proposed to alleviate the quadratic complexity of self-attention, including sparse attention [19], and linear attention [20].
However, a unified principle that jointly achieves interpretability and efficiency in self-attention remains largely unexplored. This leaves the design of efficient attention mostly heuristic and the theories behind interpretable attention less instructive. To bridge this gap, a recent work [21] introduced the the maximal coding rate reduction (MCR2[22] objective but left several important questions unresolved. Their token-statistic self-attention (TSSA) adaptively scales feature channels based on their second moments, achieving high interpretability with linear computation complexity. Nevertheless, the channel attention [23, 24] manner of TSSA contrasts sharply with the spatial manner of standard softmax attention [1]. Therefore, it remains an open question whether an interpretable framework can be developed to integrate attention mechanisms, ranging from full attention  [^1] Softmax attention calculates all possible pairwise similarities among input tokens, i.e., it models a fully-connected digraph, and and is therefore also called full attention. to linear attention and channel attention, that trade expression capacity for increased efficiency while revealing their fundamental differences.
In this paper, we propose a unified optimization objective to derive an inherently interpretable and efficient attention mechanism. To be specific, the proposed objective compresses input tokens into a union of subspaces by contracting a few representative tokens (abbreviated as representatives), making it more efficient than processing all tokens individually. By unrolling the optimization procedure over the objective, we derive our Contract-and-Broadcast Self-Attention (CBSA), which contracts a few representatives and broadcasts the contractions back to all input tokens, as illustrated in Fig. 1.
Given a fixed number of representatives, CBSA’s computational and memory complexity scales linearly with the number of tokens. Moreover, varying the choice of representatives yields CBSA variants that trade efficiency for expression capacity, or vice versa. We demonstrate that: a) when each token serves as its own representative, CBSA recovers softmax attention; b) when representatives are orthogonal (e.g., the principal directions), CBSA reduces to linear attention; c) when representatives are orthogonal but fixed (i.e., input-agnostic), CBSA degenerates into channel attention. Our findings suggest that differences among these attention mechanisms arise from distinct choices of representatives, whose number and structure induce divergent patterns of information propagation.
Figure 1: Contract-and-Broadcast Self-Attention (CBSA). Besides projection onto the subspace and back to the ambient space, our CBSA generally involves four steps: initialization, extraction, contraction, and broadcast. In Section 3.3, we show that different representative extraction strategies can simplify or fuse some of these steps, thereby leading to different attention mechanisms. A more easy to understand
Paper contributions
  1. 1.
    We propose a unified optimization objective that compresses all tokens by contracting a few representatives, thereby deriving an inherently interpretable and efficient attention mechanism.
  2. 2.
    By varying the choice of representatives, the derived mechanism, CBSA, generalizes existing attention mechanisms as special cases, reducing their fundamental differences to the number and structure of representatives.
  3. 3.
    We validate the interpretability and efficiency of CBSA through extensive experiments on visual tasks, highlighting its competitive or superior performance.
2 Notations and preliminaries
Given a positive integer $n$, we denote $n]\doteq\{1,2,\ldots,n\}$. For $s\geq t$, we use $\mathcal{O}(s,t)\subseteq\mathbb{R}^{s\times t}$ to denote the set of $s\times t$ matrices with orthonormal columns, and $\mathcal{O}(s)\doteq\mathcal{O}(s,s)$ to denote the set of $s\times s$ orthogonal matrices. We denote an identity matrix of size $n$ by $\mathbf{I}_{n}$, and similarly a zero square matrix by $\mathbf{O}_{n}$. Given a vector $\boldsymbol{v}\in\mathbb{R}^{n}$, let $\operatorname{Diag}(\boldsymbol{v})\in\mathbb{R}^{n\times n}$ be a diagonal matrix with the elements of $\boldsymbol{v}$ along the diagonal.
We use $\boldsymbol{Z}\in\mathbb{R}^{d\times N}$ to denote a sequence of $N$ input tokens represented in the $d$-dimensional ambient space, and specify these fed into the $\ell$-th attention layer as $\boldsymbol{Z}^{\ell}$. The same holds for the tokens’ representatives $\boldsymbol{Q}\in\mathbb{R}^{d\times m}$, where $m$ can be much smaller than $N$. The low-dimensional structure of representation is modeled as $K$ incoherent $p$-dimensional subspaces spanned by orthonormal bases $\boldsymbol{U}_{[K]}\doteq\{\boldsymbol{U}_{k}\in\mathcal{O}(d,p)\}_{k=1}^{K}$, such that $pK=d$ and $\boldsymbol{U}_{i}^{\top}\boldsymbol{U}_{j}=\mathbf{O}_{p},\forall i\neq j$. For the $i$-th basis vector of $\boldsymbol{U}_{k}$, we denote by $\boldsymbol{u}_{ki}$. Similar to [16], we implement $\boldsymbol{U}_{[K]}$ as learnable parameters in each attention layer, denoting the set modeled by the $\ell$-th layer as $\boldsymbol{U}^{\ell}_{[K]}=\{\boldsymbol{U}_{k}^{\ell}\}_{k=1}^{K}$.
To quantify compactness, i.e., the extent to which the tokens are compressed towards the subspaces, we adopt the (lossy) coding rate [25], which measures how efficiently the token distribution can be covered by $\epsilon$-balls under a given quantization precision $\epsilon>0$. The coding rate of the input tokens in the ambient space is defined as:
$R(\boldsymbol{Z})\doteq\frac{1}{2}\log\det\left(\mathbf{I}_{N}+\frac{d}{N\epsilon^{2}}\boldsymbol{Z}^{\top}\boldsymbol{Z}\right),$ (1)
and similarly defined for the projections onto the $k$-th subspace as $R(\boldsymbol{U}_{k}^{\top}\boldsymbol{Z})$.
3 Methods
3.1 Compressing all by contracting a few
Before presenting our proposed optimization objective, we start from a well-established learning objective grounded in compression, maximal coding rate reduction (MCR2[22, 26], and streamline it by introducing representatives.
The MCR2 objective aims to compress tokens into $K$ subspaces while expanding them in the ambient space to learn compact and structured representation. Specifically, it quantifies both compression and expansion through the coding rate, and can be formulated as follows:
$\max_{\boldsymbol{Z}}\ \Delta R(\boldsymbol{Z})\doteq\underbrace{R(\boldsymbol{Z}))}_{\text{expansion}}-\underbrace{R_{c}(\boldsymbol{Z}\mid\boldsymbol{U}_{[K]})}_{\text{compression}}\doteq R(\boldsymbol{Z})-\sum_{k=1}^{K}R(\boldsymbol{U}^{\top}_{k}\boldsymbol{Z}).$ (2)
By unrolling the optimization over the compression term of (2), Yu et al. [16] derived an inherently interpretable softmax attention mechanism called multi-head subspace self-attention (MSSA) (9), which also scales quadratically. This complexity essentially arises from the fact that (1) calculates the Gram matrix of tokens in order to compress them. [^2] By the commutative property of the coding rate [25], the Gram matrix can be equivalently replaced by the covariance matrix, which directly addresses the quadratic complexity [27, 21]. However, as discussed in the introduction, this leads to an invertible degeneration to channel attention.
Inspired by the scaling strategies built upon landmarks [28, 29, 30], in this paper, We propose to streamline the MCR2 objective by compressing all input tokens through contracting a small number of representatives of them, resulting in a new objective that requires minor modifications to the compression term of (2):
$\max_{\boldsymbol{Z}}\ R(\boldsymbol{Z})-\sum_{k=1}^{K}R(\boldsymbol{U}^{\top}_{k}\boldsymbol{Q})\quad\text{s.t.}\ \sum_{k=1}^{K}R(\boldsymbol{U}^{\top}_{k}\boldsymbol{Q})=\sum_{k=1}^{K}R(\boldsymbol{U}^{\top}_{k}\boldsymbol{Z}),$ (3)
where the optimization variable involves only input tokens $\boldsymbol{Z}$, as representatives $\boldsymbol{Q}$ is extracted from it. [^3] Note that the representatives are not necessarily a subset of the input tokens; rather, they resemble the cluster centroids in $K$-means and are continuously determined, rather than discretely selected from the inputs. When the constraint of (3) holds, (3) is equivalent to (2) but with much higher efficiency, because the number of representatives (e.g., $m=p=\nicefrac{{d}}{{K}}$) is usually far smaller.
Relaxation and regularization of (3). As the equality constraint of (3) is overly restrictive in practice, we relax it by introducing a tolerance parameter $\tau$, which bounds the absolute difference between the two coding rates within each subspace, resulting in the following inequality constraints:
$\left|R(\boldsymbol{U}^{\top}_{k}\boldsymbol{Q})-R(\boldsymbol{U}^{\top}_{k}\boldsymbol{Z})\right|\leq\tau,\ \forall k\in[K].$ (4)
As a result, contracting the representatives via (3) ensures that the input tokens are correspondingly compressed in most cases, despite a potential significant gap between their coding rates. [^4] In Fig. 6, we observe that although the gap (i.e., $\tau$) is significant, the trends of the two coding rates still closely echo each other. We thus do not deliberately constrain the value of $\tau$. Additionally, note that the coding rate is invariant under any orthogonal transformation [22]. In extreme cases, if no additional structural constraint linking representatives to input tokens is imposed, the representatives could even be orthogonal to all input tokens, while still yielding identical coding rates. Therefore, we regularize the relaxed constraints (4) through a reconstruction term, such that the representatives can be expressed as linear combinations of the input tokens.
Eventually, the optimization objective that our subsequent derivations rely on is formulated as:
$\min_{\boldsymbol{Z}}$ $\ \sum_{k=1}^{K}R(\boldsymbol{U}^{\top}_{k}\boldsymbol{Q})$
s.t. $\ \left|R(\boldsymbol{U}^{\top}_{k}\boldsymbol{Q})-R(\boldsymbol{U}^{\top}_{k}\boldsymbol{Z})\right|\leq\tau,\ \boldsymbol{U}^{\top}_{k}\boldsymbol{Q}=\boldsymbol{U}^{\top}_{k}\boldsymbol{Z}\boldsymbol{A}_{k},\ \forall k\in[K],$ (5)
where $\{\boldsymbol{A}_{k}\in\mathbb{R}^{N\times m}\}^{K}_{k=1}$ are the coefficient matrices over the dictionaries $\{\boldsymbol{U}_{k}^{\top}\boldsymbol{Z}\in\mathbb{R}^{p\times N}\}_{k=1}^{K}$. Following [16], we optimize the expansion term $R(\boldsymbol{Z})$ in alternation with the compression term via a feed-forward module named ISTA, and thus have omitted it in (5).
3.2 Contract-and-Broadcast Self-Attention
In this subsection, we derive our inherently interpretable and efficient attention mechanism by translating the gradient-based optimization procedure over (5) into forward-pass operations.
Representative initialization and extraction. According to [31], an inherently interpretable cross-attention mechanism based on MCR2, called multi-head subspace cross-attention (MSCA), can be interpreted as aligning the coding rates of its query and key/value  [^5] In the context of inherently interpretable attention mechanisms based on MCR2, the query, key, and value coincide in self-attention, while the key and value coincide in cross-attention. The reason for this will become clear in the following derivation of CBSA. Therefore, we begin with an initial guess of the representatives as the query of MSCA, and take the input tokens as the key and value to extract representatives that satisfy the constraints in (5). In addition, the attention matrices in the attention heads of MSCA  [^6] That is, $\boldsymbol{A}_{k}=\operatorname{softmax}\left((\boldsymbol{U}^{\top}_{k}\boldsymbol{Z})^{\top}(\boldsymbol{U}^{\top}_{k}\boldsymbol{Q}_{0})\right)$, where $\boldsymbol{Q}_{0}$ denotes the initialized representatives. directly serve as the coefficient matrices in (5), since the query is updated as linear combinations of the key. Following [30], we employ a pooling-based strategy for representative initialization. [^7] Another intuitive choice is to initialize them as learnable parameters, as in [32]. However, we find that this strategy suffers from a low-rank problem, which is discussed in the Appendix A. More details are deferred to the Appendix A, as our primary focus here is on the subsequent derivation.
Representative contraction and contraction broadcast. Having initialized the representatives via pooling the input tokens and extracted those satisfying the constraints of (5) via MSCA, we move on to taking a gradient descent step on the objective of (5):
$\boldsymbol{Z}\leftarrow\boldsymbol{Z}-\eta\nabla_{\boldsymbol{Z}}\left(\sum_{k=1}^{K}R(\boldsymbol{U}^{\top}_{k}\boldsymbol{Q})\right)=\boldsymbol{Z}-\kappa\operatorname{CBSA}(\boldsymbol{Z}\mid\boldsymbol{U}_{[K]}),$ (6)
[…]
where $\eta$ denotes the original step size of the gradient descent update, and the step-size $\kappa$ is introduced to absorb the gradient coefficient for notational simplicity. The contraction term in (7) is named so because it is proportional to the gradient of the representatives’ coding rate with respect to themselves, which specifies the contraction direction (abbreviated as contractions) of them. The broadcast term in (7), an $m\times N$ matrix encoding the linear relationships between representatives and input tokens, serves to broadcast the contractions back to all input tokens.
To reduce the computational overhead of the contraction step, we adopt [16, Approximation 4], which replaces the matrix inverse with a Gram matrix followed by a softmax (i.e., an attention matrix), yielding the CBSA operator implemented in our Transformer models: [^8] Note that the scaling factor and possible sign inversion introduced by this approximation are absorbed by $\kappa$ again, which is implemented as a learnable parameter in the model.
[…]
The self-attention operation above corresponds to softmax attention, except that the linear projections for query, key, and value are all replaced by the subspace basis, i.e., $\boldsymbol{W}^{\mathrm{q}}=\boldsymbol{W}^{k}=\boldsymbol{W}^{v}=\boldsymbol{U}_{k}^{\top}$. By default, we adopt this simplified version (8) in our Contract-and-Broadcast Transformer (CBT), as illustrated in Fig. 1. Nonetheless, as will be shown in the next subsection, it is the rigorous prototype (7) that exposes the fundamental connections among different attention mechanisms.
Figure 2: FLOPs of a single attention layer, with $d=384$, $H=6$, and a patch size of $16\times 16$.
Overview and complexity analysis of CBSA and CBT. Our CBSA efficiently processes input tokens through the following steps: it projects input tokens into subspaces (i.e., attention heads parameterized by $\boldsymbol{U}_{[K]}$), initializes representatives via pooling, extracts representatives via cross-attention, contracts them via self-attention, and broadcasts the contractions back to all tokens according to the attention matrices computed in the extraction step. For an overall interpretability, we adopt ISTA [16] as the feedforward block, as it optimizes the expansion term omitted from (5). We present the computational complexities of CBSA and its sub-operations in Table 1, and compare the FLOPs of different attention mechanisms in Fig. 2. Provided $N>\nicefrac{{2d}}{{H}}=2p$ (typically 128 in Transformers), the FLOPs of CBSA are lower than those of MSSA. Comparisons with more modules, such as MHSA and MLP, can be found in Fig. 9.
Table 1: Computational complexities. By default, we set $m=p=\nicefrac{{d}}{{H}}$, where $H=K$ denotes the number of attention heads (interpreted as subspaces in our case). The complexity of each sub-operation is computed by summing the costs across all heads. It is worth noting that the projection operations, which are essential to almost all attention mechanisms, bottlenecks the overall complexity always over $\mathcal{O}(Nd^{2})$.
$\mathrm{\Omega}(\operatorname{MSSA})$ $\mathrm{\Omega}(\operatorname{CBSA})$ sub-operations of CBSA
$\mathrm{\Omega}(\text{extraction})$ / $\mathrm{\Omega}(\text{broadcast})$ $\mathrm{\Omega}(\text{contraction})$ $\mathrm{\Omega}(\text{projection})$
$2Nd^{2}+2N^{2}d$ $2Nd^{2}+3Nmd+2m^{2}d$ NmdNmd $m^{2}d$ $Nd^{2}$
3.3 Variants of CBSA
In this subsection, we demonstrate that specifying different sets of representatives or their extraction strategies derives variants of CBSA that can be categorized as existing attention mechanisms. In these variants, the initialization, extraction, contraction, and broadcast steps of CBSA may not be explicitly observed, as some steps are simplified or fused together due to their specific number and structure of the representatives.
Agent Attention [30] as CBSA variant. As illustrated in Fig. 3, if we remove the contraction step of CBSA, it derives the Agent Attention, which was proposed to be a generalized form of linear attention. [^9] Since query, key, and value coincide in CBSA (8), the broadcast matrix in Agent Attention can be obtained directly from the extraction (i.e., cross-attention) step. Actually, we argue that, in this case, the contraction step is not removed but fused into the extraction step, since most of the extraction’s purpose already achieved by pooling the input tokens.
We previously extract representatives satisfying the relaxed constraints in (5) via MSCA with an initial guess, yielding an approximate numerical solution to the constraints of the original problem (3). Beyond this, closed-form optimal solutions  [^10] That is, the representatives that not only have the same coding rate as the input tokens but also possess an explicit coefficient matrix allowing them to be linearly expressed by the input tokens. also exist, which can either improve interpretability and efficiency or enhance the expressive capacity of the resulting mechanism.
Softmax attention as CBSA variant. Apparently, the input tokens themselves satisfy the constraints and thus can be directly used as the representatives, i.e., $\boldsymbol{Q}=\boldsymbol{Z},m=N$. In this case, we say the input tokens are self-expressed [33], since data samples are linearly represented over a dictionary composed of themselves. Then we can take a trivial solution for the regularization term where all coefficient matrices reduces to identity matrices. Substituting them into (8) yields the following operator, known as MSSA in white-box transformers [16]:
[…]
Linear attention as CBSA variant. To analyze the case of orthogonal representatives, we start with the canonical choice: the principal directions of input tokens, along which projections yield the principal components. Specifically, We perform singular value decomposition (SVD) on the token projection within each subspace:
$\boldsymbol{U}^{\top}_{k}\boldsymbol{Z}=\boldsymbol{L}_{k}\boldsymbol{\Sigma}_{k}\boldsymbol{R}_{k}^{\top},\ \ \forall k\in[K],$ (10)
where $\boldsymbol{L}_{k}\in\mathcal{O}(p)$, $\boldsymbol{R}_{k}\in\mathcal{O}(N,p)$, and $\boldsymbol{\Sigma}_{k}$ is a $p\times p$ diagonal matrix of singular values. The columns of $\boldsymbol{L}_{k}$ are known as the principal directions. Then, by right multiplying both sides by $\boldsymbol{R}_{k}$, we have:
$\boldsymbol{U}^{\top}_{k}\boldsymbol{Z}\boldsymbol{R}_{k}=\boldsymbol{L}_{k}\boldsymbol{\Sigma}_{k}\boldsymbol{R}_{k}^{\top}\boldsymbol{R}_{k}=\boldsymbol{L}_{k}\boldsymbol{\Sigma}_{k},\ \ \forall k\in[K],$ (11)
By comparing (10) with the constraints in (5), we let $\boldsymbol{U}_{k}^{\top}\boldsymbol{Q}=\boldsymbol{L}_{k}\boldsymbol{\Sigma}_{k},\boldsymbol{A}_{k}=\boldsymbol{R}_{k}$. Substituting them into (7) leads to the following operation:
[…]
where $F$ is a function defined on the spectrum of a positive semi-definite matrix which applies $f(\lambda_{i})=\nicefrac{{\epsilon^{2}}}{{(\epsilon^{2}+\lambda_{i})}}$ to the eigenvalues $\{\lambda_{i}\}_{i=1}^{p}$ of the covariance matrix of the token projections. [^11] As $f$ is monotonically decreasing, (12) preserves the principle directions (i.e., representatives) with large variance while suppressing those with vanishing variance [27]. In Appendix B, we prove that this result holds for any set of orthogonal representatives. We find that (12) conceptually resembles linear attention, as it also factorizes the $N\times N$ attention matrix and multiplies the key and value first to linearize the computational complexity. [^12] In fact, linear attention is more expressive than (12), since $(\boldsymbol{U}^{\top}_{k}\boldsymbol{Z})(\boldsymbol{U}^{\top}_{k}\boldsymbol{Z})^{\top}$ is symmetric, whereas $(\boldsymbol{W}^{v}\boldsymbol{Z})(\boldsymbol{W}^{k}\boldsymbol{Z})^{\top}$ is generally not. How to interpret the transition from such symmetric attention to their asymmetric counterparts remains largely unexplored. For now, one can view it as overparameterization.
Channel attention as CBSA variant. If we assume that the subspace bases $\boldsymbol{U}_{k}$ always correspond to the principal directions of any input token set, which is practically impossible but simplifies the operations, the representatives can be fixed as $\boldsymbol{U}_{k}$ (i.e., $\boldsymbol{Q}=[\boldsymbol{U}_{1}^{\top}\ldots\boldsymbol{U}_{K}^{\top}]^{\top}$), and are thus orthogonal and input-agnostic. Then (12) is reduced to:
$\sum_{k=1}^{K}\boldsymbol{U}_{k}\boldsymbol{D}_{k}\boldsymbol{U}_{k}^{\top}\boldsymbol{Z},\ \boldsymbol{D}_{k}\doteq\operatorname{Diag}\left([f\big((\boldsymbol{u}_{ki}^{\top}\boldsymbol{Z})(\boldsymbol{u}_{ki}^{\top}\boldsymbol{Z})^{\top}\big)]_{i=1}^{p}\right),$ (13)
which corresponds to TSSA [21] without exploiting token-to-subspace membership. In (13), the feature channels are adaptively scaled based on the second moment of the token projections, whereas typical channel attention mechanisms employ an MLP to predict the scaling factors.
Besides a graph perspective, the above findings can also be interpreted from a dictionary learning perspective, by treating the representatives as atoms of a dictionary. When the representatives are orthogonal and fixed, they form a complete dictionary; when they are orthogonal but input-dependent, they resemble a submatrix of an overcomplete dictionary in compressed sensing [34]. When the input tokens themselves serve as representatives, they constitute a self-expressive dictionary [33].
Figure 3: Overview of different attention mechanisms. Zoom in for better view.
4 Experiments
In this section, we evaluate the interpretability and efficiency of the proposed inherently interpretable attention, CBSA, and the CBT models built upon it. Since natural images often lie on low-dimensional subspaces [35, 36], we focus on classical visual tasks such as image classification and semantic segmentation, where higher resolutions generally lead to better accuracy [37, 38]. Experimental results copied from their original work will be in grey color, otherwise the results will be reproduced by us in varied settings for fair comparisons.
Baseline and training configuration. We primarily compare CBSA with previous interpretable attention mechanisms grounded in MCR2, along with pioneering and classical Transformer-based models. Table 2 summarizes these baselines with brief descriptions. Results directly copied from the original papers are shown in grey, while the others are reproduced by us under varied settings for fair comparisons. Unless otherwise specified, the training configuration follows that of the baselines, with detailed information provided in Appendix C.
Implementation detail. We over-parameterize the projection step that maps tokens back to the ambient space, where the projection matrix should theoretically be $\boldsymbol{U}_{[K]}$, by a set of independently learned parameter matrices. This trick is also adopted by MSSA [16] and TSSA [21], and its impact has been investigated in [39]. In short, although this relaxation compromises the theoretical rigor, it is crucial for achieving better accuracy. In addition, the positive scalar $\kappa$ in (6) is implemented as a learnable parameter without constraining its sign. This allows the model to flexibly choose between compression and de-compression. We provide the PyTorch implementation in Appendix D.
Table 2: Summary of baselines. Note that these methods are not necessarily limited to the task listed behind, our description here only reflects their usages in this paper.
Method Attention mechanism Complexity Interpretable Task
ViT [2] MHSA (softmax attention) quadratic image classification
CRATE [16] MSSA (softmax attention) quadratic image classification
ToST [21] TSSA (channel attention) linear image classification
Agent Attention [30] Agent Attention (linear attention) linear image classification
Segmenter [40] MHSA quadratic semantic segmentation
DEPICT [31] MSSA / MSCA (cross-attention) quadratic / linear semantic segmentation
4.1 Advantages enabled by interpretability
In this subsection, we show some superior advantages of CBSA over non-interpretable attention mechanisms. The most essential one is its interpretability, which further induces other desirable properties such as robustness and emergent segmentation.
(a) Raw
(b) 1 iters
(c) 256 iters
(d) 512 iters
(e) 640 iters
(f) 768 iters
(g) 896 iters
(h) 1024 iters
Figure 4: Compact and structured representation. There are 10 classes in total and each class has 200 samples.
Learning compact and structured representation. Similar to MCR2, our optimization objective (3) also aims to learn compact and structured representations, i.e., compressing input tokens towards low-dimensional structures such as a mixture of subspaces. To verify that CBSA can achieve this goal, we evaluate the linear-attention-like variant of CBSA (12) on synthetic data. As illustrated in Fig. 4, we emulate image tokens as points in a 3-dimensional space, where ground-truth clustering memberships are indicated by color, and apply (12) iteratively. Since the operation is directly performed in the ambient space and no parameterized subspace is required, this procedure can be executed in a purely forward manner under supervision.
Figure 5: We measure the compression term of (2) across layers in different models.
Compressing all by contracting a few. To verify our interpretation that CBSA compresses all tokens through the contraction of a few representatives, we conduct two-fold experiments showing that the input tokens are indeed compressed and that this compression is driven by the contraction of the representatives. In Fig. 5, we measure the compression term of (2) using the coding rate (1) and its normalized variant  [^13] All input tokens are first normalized to unit vectors before calculating the normalized coding rate.. The normalized coding rate is invariant to in-place scaling and therefore captures only the angular compression. We observe two pieces of evidence supporting the compression interpretation: 1) representation compactness in the last layer, measured by the coding rate in Fig. 5, positively correlates with classification accuracy in Table 3; 2) in nearly all models, the latter half of the layers exhibit consecutive compression. We also identify a de-compression phenomenon in the earlier layers, which we regard as a “known unknown” requiring further investigation. In Fig. 6, we measure the “compressed volume” of input tokens and their representatives separately after being processed by CBSA (without projecting back to the ambient space, so the results are reported per subspace). We observe that the compressed volumes of the two exhibit highly similar trends across most subspaces.
Figure 6: We measure the compressed volume of input tokens and their representatives using the reduced coding rate across all heads of a model. Results from CBT-T/S are presented here, those of CBT-B/L can be found in Fig. 10. Note that we set $\kappa=1$ to exclude the de-compression cases observed in Fig. 5.
Figure 7: [CLS] attention map. We use $\boldsymbol{A}_{k}\boldsymbol{A}_{k}^{\top}\in\mathbb{R}^{N\times N}$ to empirically approximate the full attention matrix. Based on this, we extract the attention map of the [CLS] token from the early, middle, and late layers of CBT, CRATE, and their hybrid model.
Emergent segmentation properties. Owing to MSSA, segmentation properties emerge in CRATE with merely standard supervised classification training [41]. In Fig. 7, we visualize the attention map of the [CLS] token across different models. We observe that CBT attends to more semantically meaningful regions in the early layers compared to CRATE, but its segmentation properties do not persist in subsequent layers. To address this, we construct a hybrid model, termed CRATE+CBT, where the first half of the attention layers employ MSSA and the latter half employ CBSA. In this hybrid model, segmentation properties not only emerge in the very first layer, but are also progressively enhanced in the following layers, rather than fading as in CBT. Qualitative results supporting this conclusion can be found in Appendix C.
Figure 8: Semantic segmentation on ADE20K. $\dagger$ indicates that CBSA implemented rigorously, i.e., without the overparameterization trick. All results above are evaluated on the ADE20K validation set. Leftmost panel: The dotted horizontal line marks the best result. Right two panels: Random Gaussian noise with zero mean and standard deviation $\sigma$ is independently added to each parameter of the attention layers in the decoder.
Robustness under parameter perturbation. As CBSA models low-dimensional subspaces via their attention heads, with parameter matrices corresponding to subspace bases, perturbing these basis vectors with relatively small noise does not significantly alter the subspaces they span [31]. Consequently, CBSA exhibits extreme robustness to parameter perturbations, whereas the black-box method Segmenter [40] collapses under the same perturbations, as illustrated in the right two panels of Fig. 8.
Table 3: Image classification on ImageNet-1K. The models are trained on images of resolution $224\times 224$ with a patch size of $16$. Note that the CBT models and ViT employ convolutional embedding layers, whereas the CRATE models use linear embeddings.
Datasets CBT-T(iny) CBT-S(mall) CBT-B(ase) CBT-L(arge) CRATE-B CRATE-L ViT-S
# parameters 1.8M 6.7M 25.7M 83.1M 22.8M 77.6M 22.1M
FLOPs 1.1G 4.0G 15.1G 47.3G 12.6G 43.3G 9.8G
ImageNet-1K 63.2 71.4 73.4 74.4 70.8 71.3 72.4
CIFAR10 94.8 96.3 96.7 97.3 96.8 97.2 97.2
CIFAR100 76.5 80.4 82.0 83.4 82.7 83.6 83.2
Oxford Flowers-102 88.4 91.7 93.6 93.9 88.7 88.3 88.5
Oxford-IIIT-Pets 86.8 91.6 92.6 92.9 85.3 87.4 88.6
Table 4: Fair comparisons on ImageNet-1K. All models employ convolutional embedding layers and ISTA feedforward blocks. Their only difference lies in the attention mechanism. Specially, the Agent models do not follow the original implementation in [30]; instead, they are implemented as CBSA without the contraction step, as derived in Section $3.$
ImageNet-1K CBSA-T CBSA-S TSSA-T TSSA-S MSSA-T MSSA-S Agent-T Agent-S
# pairwise similarities 0.53M 1.1M 0.45M 0.91M 1.4M 2.8M 0.52M 1.0M
Top 1 accuracy 63.2 71.4 61.2 68.5 64.7 72.1 63.8 71.8
Figure 9: Adapting pre-trained ViTs into CBSA style. We finetune both the adapted models and the original ViTs on ImageNet-1K for 50 epochs, and report the top-1 accuracy on the validation set. CBSA leverages the three distinct projection matrices inherited from the pretrained MHSA to calculate query, key, and value, instead of using a single projection matrix as in (8). CBSA refers to CBSA without the contraction step, which essentially becomes Agent Attention.
4.2 Evaluations on real-world visual tasks
We pretrain CBT models on the ImageNet-1K dataset, and finetune them on several downstream datasets. The top-1 accuracy on the validation sets is reported in Table 3. In particular, our CBT-Small achieves comparable top-1 accuracy to ViT-S using only 30% of the parameters and 40% of the FLOPs. Compared to CRATE models built upon [16], our CBT models (with convolutional embedding layers) perform remarkably better while using fewer parameters and FLOPs.
We also conduct fair comparisons across different attention mechanisms, which can be regarded as variants of CBSA, and report the results as in Table 4. In this setting, our CBT models remain competitive with CRATE while computing significantly fewer pairwise similarities. These results confirm that, from MSSA to CBSA and then to TSSA, the number of pairwise similarity computations decreases at the cost of some performance. Throughput comparisons on high-resolution images ($512\times 512$) are presented in Table 7, showing that our method consistently achieves superior training and inference efficiency.
To investigate the potential of applying CBSA to large-scale pretraining, we preliminarily finetune ViT models pretrained on ImageNet-21K [^14] The checkpoint is obtained from the timm [42] library. by adapting their attention blocks into the CBSA style. For comparison, we also adapt them into linear attention, with results shown in Fig. 9. Although CBSA deviates more from MHSA than linear attention and is consequently harder to adapt from pretrained ViTs, CBSA achieves performance comparable to linear attention with nearly the same FLOPs. Moreover, when adapted to Agent Attention, CBSA surpasses linear attention, which has been carefully experimented in [30].
For semantic segmentation, we build CBT decoders by stacking CBSA layers without feed-forward modules on top of ViT encoders, following the design of [31]. We evaluate them on the ADE20K dataset [43], with results shown in the left panel of Fig. 8. Our CBT decoder consistently surpasses both white-box (DEPICT) and black-box (Segmenter) counterparts that rely on softmax attention. In particular, the best-performing CBT decoder improves upon Segmenter by $1.5\%$ mIoU while using only $20\%$ of the FLOPs and $0.06\%$ of the pairwise similarities in the decoder.
5 Related work
Efficient attention mechanisms can be roughly divided into two categories: sparse attention and linear attention [44]. Sparse attention approaches sparsify the attention matrix proactively by restricting the attention span to random, fixed [45, 19], learnable [46, 47] patterns, or their combinations. Linear attention [20, 48] decomposes the attention matrix into a product of two low-rank matrices and thus avoids its explicit computation via the associative property of multiplication. The idea of using representative tokens has been applied to both. Global tokens or memory can be introduced in sparse attention to maintain global connectivity, thereby further shrinking the attention span [49, 50]. Meanwhile, various works are implicitly sparse for reducing the number of input tokens [51, 52, 53, 54]. Agent tokens [30] and landmarks [28] can also be incorporated into linear attention from different perspectives.
Our CBSA distinguishes itself from previous efficient attention mechanisms by being inherently interpretable and derived from an optimization objective that efficiently compresses input tokens toward low-dimensional structures. Moreover, it can not only be viewed as sparse attention or linear attention through the representatives, but also mathematically generalizes linear attention and channel attention as special cases.
6 Conclusion
We proposed an optimization objective that compresses all input tokens by contracting a small number of representatives, thereby unifying the investigation towards interpretability and efficiency. By unrolling the gradient optimization steps for this objective, we derived an inherently interpretable and efficient attention mechanism, called CBSA. We found that many existing attention mechanisms can be generalized within the CBSA framework by specifying the number and structure of representatives, thus revealing their fundamental connections. We further validated the effectiveness of CBSA through extensive experiments on visual tasks. Therefore, we demonstrated that the preliminary framework established in this work offers a promising direction for exploring new attention mechanisms in an interpretable and efficient way.
Appendix
Appendix A Representative initialization
Besides pooling the input tokens, another intuitive choice for representative initialization is to set them as learnable (i.e., trainable) parameters, similar to the object queries in [32], the class embeddings in [40], and the registers in [55]. However, in this case we found that the attention matrix in the extraction (cross-attention) step [^15] That is, $\operatorname{rank}\left(\operatorname{softmax}\left((\boldsymbol{U}_{k}^{\top}\boldsymbol{Z})^{\top}(\boldsymbol{U}_{k}\boldsymbol{Q})\right)\right)$. is extremely low rank. For example, although its maximal rank can be $m=p=\frac{d}{H}=64$ in CBT models, it typically remains around $5$. This indicates that far fewer initialized representatives are being utilized than expected, leading to unsatisfactory performance of CBSA.
By contrast, when pooling-based initialization is employed, the rank of the attention matrix always attains its maximum, i.e., $m$. Nonetheless, as pointed out in [56], this still constitutes a low-rank bottleneck: linear attention underperforms softmax attention, whose attention matrix has rank $N\gg m$. Although this issue can be alleviated by introducing depth-wise convolution (DWC), which boosts performance while preserving linear complexity, our work does not involve this modification since we focus on theoretical contributions rather than empirical improvements.
Appendix B Derivations
Here, we show how (12) is derived by substituting $\boldsymbol{U}_{k}^{\top}\boldsymbol{Q}=\boldsymbol{L}_{k}\boldsymbol{\Sigma}_{k},\boldsymbol{A}_{k}=\boldsymbol{R}_{k}$ into (7):
$\sum_{k=1}^{K}\boldsymbol{U}_{k}\boldsymbol{L}_{k}\boldsymbol{\Sigma}_{k}\left(\mathbf{I}_{m}+\frac{m}{m\epsilon^{2}}(\boldsymbol{L}_{k}\boldsymbol{\Sigma}_{k})^{\top}(\boldsymbol{L}_{k}\boldsymbol{\Sigma}_{k})\right)^{-1}\boldsymbol{R}_{k}^{\top}$ (14)
$=$ $\sum_{k=1}^{K}\boldsymbol{U}_{k}\boldsymbol{L}_{k}\boldsymbol{\Sigma}_{k}\left(\mathbf{I}_{m}+\frac{1}{\epsilon^{2}}\boldsymbol{\Sigma}_{k}^{\top}\boldsymbol{L}_{k}^{\top}\boldsymbol{L}_{k}\boldsymbol{\Sigma}_{k}\right)^{-1}\boldsymbol{R}_{k}^{\top}$ (15)
$=$ $\sum_{k=1}^{K}\boldsymbol{U}_{k}\boldsymbol{L}_{k}\boldsymbol{\Sigma}_{k}\left(\mathbf{I}_{m}+\frac{1}{\epsilon^{2}}\boldsymbol{\Sigma}_{k}^{2}\right)^{-1}\boldsymbol{R}_{k}^{\top}$ (16)
$=$ $\sum_{k=1}^{K}\boldsymbol{U}_{k}\boldsymbol{L}_{k}\left(\mathbf{I}_{m}+\frac{1}{\epsilon^{2}}\boldsymbol{\Sigma}_{k}^{2}\right)^{-1}\boldsymbol{\Sigma}_{k}\boldsymbol{R}_{k}^{\top}$ (17)
$=$ $\sum_{k=1}^{K}\boldsymbol{U}_{k}\boldsymbol{L}_{k}\left(\mathbf{I}_{m}+\frac{1}{\epsilon^{2}}\boldsymbol{\Sigma}_{k}^{2}\right)^{-1}\boldsymbol{L}_{k}^{\top}\boldsymbol{U}_{k}^{\top}\boldsymbol{Z}$ (18)
$=$ $\sum_{k=1}^{K}\boldsymbol{U}_{k}F\left(\boldsymbol{L}_{k}\boldsymbol{\Sigma}_{k}^{2}\boldsymbol{L}_{k}^{\top}\right)\boldsymbol{U}_{k}^{\top}\boldsymbol{Z}$ (19)
$=$ $\sum_{k=1}^{K}\boldsymbol{U}_{k}F\left((\boldsymbol{U}_{k}^{\top}\boldsymbol{Z})(\boldsymbol{U}_{k}^{\top}\boldsymbol{Z})^{\top}\right)\boldsymbol{U}_{k}^{\top}\boldsymbol{Z}$ (20)
Appendix C More experimental results
Training setup. We train the CBT models in Table 3 and all models in Table 4 150 epochs with the Lion optimizer [57]. The learning rate is $2.0\times 10^{-4}$, the weight decay coefficient is $0.05$, and the batch size is $256$. We also incorporate a warm-up strategy over the first $20$ epochs. For data augmentation, we adopt a rather simple choice: just random cropping and random horizontal flipping. We apply label smoothing with a smoothing coefficient of $0.1$. For fine-tuning, we use the AdamW optimizer [58], a learning rate of $5\times 10^{-5}$, weight decay of 0.010.0-1, and batch size $64$. The settings above are largely inherited from [16].
Quantitative segmentation properties. Following prior works [7, 41], we evaluate the zero-shot segmentation performance of ImageNet-1K pretrained models on the PASCAL VOC12 validation set [59]. Specifically, we assess the best-performing attention maps of the [CLS] token, but in a simplified setting that considers only three segmentation targets, instead of fine-grained semantic classes. In Table 5, we find that the hybrid model has consistently better segmentation performance. We also measure the segmentation performance of different layers in Table 6.
Table 5: Zero-shot segmentation. We use the jaccard similarity, which is also called intersection over union (IoU), to quantify the alignment between the [CLS] token’s attention map and the segmentation ground truth. The best-performing results are highlighted in bold.
Segmentation target (CRATE+CBT)-S CRATE-S CBT-S ToST-S
Foreground 0.68 0.65 0.51 0.42
Background 0.78 0.76 0.72 0.75
Boundary 0.18 0.17 0.15 0.12
Table 6: Zero-shot segmentation across layers. The top-performing three layers for each model are underlined.
Models L11 L22 L33 L44 L55 L66 L77 L88 L99 L1010 L1111 L1212
(CRATE+CBT)-S 0.37 0.44 0.55 0.46 0.56 0.58 0.60 0.53 0.59 0.56 0.55 0.55
CRATE-S 0.34 0.45 0.51 0.48 0.53 0.56 0.53 0.57 0.55 0.52 0.50 0.39
CBT-S 0.39 0.41 0.43 0.42 0.37 0.40 0.38 0.33 0.35 0.34 0.36 0.33
Table 7: Throughput comparisons. The results below are experimented on the CIFAR-10 dataset with an image resolution of $512\times 512$.
im/sec CBT-T ViT-T CRATE-T ToST-T CBT-S ViT-S CRATE-S ToST-S
training 336 174 203 323 205 94 111 199
inference 572 370 395 533 405 211 220 429
Figure 10: Additional results on CBT-B and CBT-L for the experiment in Fig. 6.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 11: Gram matrices of basis vectors modeled by the layers of CBT-S. According to our assumption in Section 2, these matrices should approximate identity matrices.
Appendix D PyTorch implementation
1
1class CBSA($n$.Module):
2 def __init__(self, dim, heads, dim_head):
3 super().__init__()
4 inner_dim = heads * dim_head
5 self.heads = heads
6 self.dim_head = dim_head
7 self.scale = dim_head ** -0.5
8 self.attend = $n$.Softmax(dim=-1)
9 self.pool = $n$.AdaptiveAvgPool2d(output_size=(8, 8))
10 # subspace bases
11 self.proj = $n$.Linear(dim, inner_dim, bias=False)
12 # over-parameterization
13 self.to_out = $n$.Linear(inner_dim, dim)
14 # step-sizes
15 self.ss_x = $n$.Parameter(torch.randn(heads, 1, 1))
16 self.ss_rep = $n$.Parameter(torch.randn(heads, 1, 1))
17
18 def attention(self, query, key, value):
19 dots = (query @ key.transpose(-1, -2)) * self.scale
20 attn = self.attend(dots)
21 out = attn @ value
22 return out, attn
23
24 def forward(self, x):
25 b, n, c = x.shape
26 height = width = int(n ** 0.5)
27 # projection onto subspaces
28 w = self.proj(x)
29 # representative initialization
30 rep = self.pool(w[:, :-1, :].reshape(b, height, width, c).permute(0, 3, 1, 2)).reshape(b, c, -1).permute(0, 2, 1)
31 w = w.reshape(b, n, self.heads, self.dim_head).permute(0, 2, 1, 3)
32 rep = rep.reshape(b, 64, self.heads, self.dim_head).permute(0, 2, 1, 3)
33 # representative extraction
34 rep_delta, attn = self.attention(rep, w, w)
35 rep = rep + self.ss_rep * rep_delta
36 # representative contraction
37 x_delta, _ = self.attention(rep, rep, rep)
38 # contraction broadcast
39 x_delta = attn.transpose(-1, -2) @ x_delta
40 x_delta = self.ss_x * x_delta
41 x_delta = rearrange(x_delta, ’b h n k -> b n (h k)’)
42 # projection back to the ambient space
43 return self.to_out(x_delta)
Algorithm 1 PyTorch implementation of CBSA (8)
1
1class CBSA($n$.Module):
2 def __init__(self, dim, heads, dim_head):
3 super().__init__()
4 inner_dim = heads * dim_head
5 self.heads = heads
6 self.dim_head = dim_head
7 self.scale = dim_head ** -0.5
8 self.attend = $n$.Softmax(dim=-1)
9 # subspace bases
10 self.proj = $n$.Parameter(self.scale * torch.randn(inner_dim, dim))
11 # over-parameterization
12 self.to_out = $n$.Linear(inner_dim, dim)
13 # step-sizes
14 self.ss_x = $n$.Parameter(torch.randn(heads, 1, 1))
15 self.ss_rep = $n$.Parameter(torch.randn(heads, 1, 1))
16
17 def attention(self, query, key, value):
18 dots = (query @ key.transpose(-1, -2)) * self.scale
19 attn = self.attend(dots)
20 out = attn @ value
21 return out, attn
22
23 def forward(self, x):
24 b, n, c = x.shape
25 # normalize the basis vectors
26 proj = self.proj / self.proj.norm(dim=-1, keepdim=True)
27 # projection onto subspaces
28 w = rearrange(x @ proj.t(), ’b n (h k) -> b h n k’, h=self.heads)
29 # representative initialization (couple them with the basis vectors but preserved the norm)
30 rep = rearrange(self.proj, ’(h k) d -> h k d’, h=self.heads)
31 rep = rep @ rearrange(proj, ’(h k) d -> h k d’, h=self.heads).transpose(-1, -2)
32 rep = rep.expand(b, -1, -1, -1)
33 # representative extraction
34 rep_delta, attn = self.attention(rep, w, w)
35 rep = rep + self.ss_rep * rep_delta
36 # representative contraction
37 x_delta, _ = self.attention(rep, rep, rep)
38 # contraction broadcast
39 x_delta = attn.transpose(-1, -2) @ x_delta
40 x_delta = self.ss_x * x_delta
41 x_delta = rearrange(x_delta, ’b h n k -> b n (h k)’)
42 # projection back to the ambient space
43 return self.to_out(x_delta)
Algorithm 2 PyTorch implementation of CBSA in Fig. 8
1
1class CBSA($n$.Module):
2 def __init__(
3 self,
4 dim: int,
5 num_heads: int = 8,
6 qkv_bias: bool = False,
7 proj_bias: bool = True,
8 ) -> None:
9 super().__init__()
10 assert dim % num_heads == 0, ’dim should be divisible by num_heads’
11 self.num_heads = num_heads
12 self.head_dim = dim // num_heads
13 self.scale = self.head_dim ** -0.5
14 self.pool = $n$.AdaptiveAvgPool2d(output_size=(8, 8))
15 self.qkv = $n$.Linear(dim, dim * 3, bias=qkv_bias)
16 self.proj = $n$.Linear(dim, dim, bias=proj_bias)
17
18 def attention(self, query, key, value):
19 query = query * self.scale
20 attn = query @ key.transpose(-2, -1)
21 attn = attn.softmax(dim=-1)
22 out = attn @ value
23 return out, attn
24
25 def forward(self, X: torch.Tensor) -> torch.Tensor:
26 B, N, C = X.shape
27 X_qkv = self.qkv(X).reshape(B, N, 3, C).permute(2, 0, 1, 3)
28 X_q, X_k, X_v = X_qkv.unbind(0)
29 H = W = int(N ** 0.5)
30 Q = self.pool(X_q[:, :-1, :].reshape(B, H, W, C).permute(0, 3, 1, 2)).reshape(B, C, -1).permute(0, 2, 1)
31 X_q = X_q.reshape(B, N, self.num_heads, self.head_dim).permute(0, 2, 1, 3)
32 X_k = X_k.reshape(B, N, self.num_heads, self.head_dim).permute(0, 2, 1, 3)
33 X_v = X_v.reshape(B, N, self.num_heads, self.head_dim).permute(0, 2, 1, 3)
34 Q = Q.reshape(B, 64, self.num_heads, self.head_dim).permute(0, 2, 1, 3)
35 Q_delta, _ = self.attention(Q, X_k, X_v)
36 Q = Q + Q_delta
37 X_delta, _ = self.attention(Q, Q, Q)
38 attn = ((X_q * self.scale) @ Q.transpose(-1, -2)).softmax(dim=-1)
39 X = attn @ X_delta
40 X = X.transpose(1, 2).reshape(B, N, C)
41 X = self.proj(X)
42 return X
Algorithm 3 PyTorch implementation of CBSA in Fig. 9